# Patterns of Patterns III

### v1

@@ 6.1. Limitations and Lessons Learned
While the outcomes of our PLACARD-based workshops (both in-person and simulated) were promising, several limitations emerged that highlight areas for refinement—both in the method itself and in our implementation of it. Generally, these limitations correspond to lessons that could be applied in future work.
@@
Cognitive demands of CLA: The abstraction and depth of Causal Layered Analysis (CLA) often overwhelmed participants, especially when scaffolding was limited. Provide lightweight onboarding tools—such as mini-patterns, worked examples, or scenario walkthroughs—to ease entry into layered analysis.
Short workshop windows (60–90 mins): Limited time forced trade-offs, with later phases (e.g., PAR) often rushed or skipped. Recognize Timebox the Core as a necessary pattern—but pair it with mechanisms for asynchronous reflection or follow-up synthesis.
@@
Sequential framing constrained emergent inquiry: Participants sometimes forced their intuitive insights into retroactive justifications. Allow more modular, remixable process designs. Embed “Improvise to Test” and “Pattern the Play” to legitimize non-linear sense-making.
Deliberative bias limits affective and tacit knowledge: Structured methods sometimes crowded out emotional or disruptive forms of insight. Integrate methods or prompts that scaffold expressive, intuitive, and affective contributions—e.g., through metaphor, narrative, or art-based reflection.
Risk of premature formalism: Design patterns encouraged generalization at the expense of context specificity. Emphasize pattern annotations and situated stories. Use pattern revision cycles to adapt and diversify rather than standardize.
Logistical overhead of documentation: Without dedicated roles or tooling, PLACARD’s emphasis on shared surfaces was burdensome. Support tool development and role formalization (e.g., for pattern capture). Design light-touch alternatives for low-resource contexts.
Simulated agents were archetypal, not adaptive: The multi-agent models lacked real-time learning or memory; they played out scripted reasoning paths. Treat the current set of simulations as design probes rather than predictive models. Use them to explore design space and pattern viability, not behavioral realism.
@@
No persistent identity or memory across sessions: Agents did not evolve across workshops, limiting insight into long-term learning. Incorporate persistent memory structures and episodic traces to study cumulative reflection, pattern evolution, and institutionalization.
Language model constraints shaped outcomes: Agent behaviors emerged from a coherence-biased generative system, not open-ended cognition. Simulations must be interpreted with epistemic humility. Their role is speculative, not definitive—ideal for surfacing design tensions, not modeling reality.
@@ 6.2. Related Work
Looking forward, the next phase of PLACARD’s evolution could be a platform for cultivating design-pattern-capable agents, communities, and learning environments. PLACARD should be understood not as a finished method but as a generative infrastructure: a living commons of epistemic tools, shaped by their user communities.  These insights can be combined with Elinor Ostrom’s Design Principles for Governing a Commons to serve as a lens for evaluating and extending PLACARD, beyond the “agent” level to the “environment” level. On this point, we can point to some work: Pitt et al. (2012) explore how institutions can self-organize by formalizing socio-economic principles that guide agent behavior within Multi-Agent Systems (MAS). Drawing inspiration from Elinor Ostrom’s work, they propose an axiomatic design approach that captures principles such as proportionality, accountability, and collective choice in a machine-interpretable form. Where Pitt et al. emphasize axioms as foundational building blocks, our PLACARD-based approach leverages design patterns as intermediate representations that bridge human practice and machine reasoning.  In a more recent paper, Mertzani and Pitt (2024) present a methodological framework for integrating power sensitivity into the design of cyber-physical systems by drawing on second- and third-order cybernetics. They propose pattern-based design methods that foreground the observer’s position, across different layes of embeddedness—with beneficial applications of technology ranging from “Saves Physical and Cognitive Resources” to “Solves Complex Problems” to “Deals with Social Constructions” to “Mitigates Existential Crises”.
@@ 7. Conclusion
Looking back over the limitations and lessons learned, it’s worth drawing out one top-level remork. ChatGPT is adept at text processing, and we have put it to use significantly in our work on this paper. It was able to generate and discuss design patterns that would describe pattern-capable agents. However, it was not able to actually use these patterns to build and run pattern-capable agents. We hope others will be interested in using LLMs to work with design patterns—but we also hope that such usage will serve as a landmark rather than a destination.
Including a computational use case within the paper suggested a key amendment to PLACARD: adding a fourth environment dimension to the pattern catalogue . Human-facing and Human-Computer Interface-facing patterns should eventually be extended similarly.
For now, to close this paper, we put forward the following CLA, which brings together the “human” and “AI” use cases and outlines a trajectory of future work. In the context of this paper, the CLA layers can aligned with the expanded view of PLACARD that we have discussed—encompassing sensory (litany), cognitive (system), motor (worldview), and environment (myth) dimensions.
@@
These are the most visible and widely discussed elements of current AI discourse: the promise and perils of large language models (LLMs) and the early success of LLM-associated code-writing agents.
@@
To move beyond isolated capabilities, we introduce pattern-based epistemic scaffolding for multi-agent collaboration. This makes the system reflexive, composable, and governed by design principles.
@@
Agents trained to recognize and apply design patterns (from the PLACARD catalogue) in their interactions. They exhibit situated competence and can engage in structured reasoning, scenario exploration, and inquiry.
@@
These agents do not just use patterns—they also monitor, revise, and propose new ones, participating in iterative design and reflective cycles. They enable epistemic learning at the system level, coordinating multi-agent behavior across contexts.
@@
Inspired by Elinor Ostrom’s principles for governing commons, we treat the MAS as a social-technological institution that evolves through interaction.
@@
This layer operationalizes developmental patterns that support reflexivity, feedback, nested structures, and shared norm evolution across MAS environments. These agents co-design systems of engagement, monitor norm congruence, and resolve governance disputes—just as human institutions do. They instantiate a worldview in which intelligence is not just personal or cognitive, but systemic and collective.
@@
Beyond technical systems, our proposal invites reflection on how real-world institutions—academic, policy, scientific—can themselves become more adaptive, transparent, and pattern-literate through interaction with these tools.
@@
This is the last layer because it is grounded in recursive reform. Humans build tools to simulate and improve human institutions—not to replace them, but to co-evolve. This layer anchors the entire architecture in civic legitimacy, ethical accountability, and reflective participation. Our claim is not merely that we can control the system, but that we can use it to help re-pattern the world we inhabit. By working with agents that are both pattern- and institution-capable, we gain assurances not only about avoiding “run-away AI” scenarios, but about avoiding other tragedies of the commons.
@@ 8. Acknowledgements
We wish to acknowledge discussions with Cris Viseu around our 2025 EuroPLoP workshop, which continues the line of work described in the paper.
This paper also benefited from extensive interaction with OpenAI’s ChatGPT. Its capacities for text generation, summarisation, pattern elicitation, and recursive prompting played a formative role in shaping both the structure and substance of this work. We also gratefully acknowledge the unnamed authors of the texts on which ChatGPT was trained. Responsibility for the interpretations, appropriations, and conclusions presented here remains entirely with the human authors.
@@
This appendix supports the analysis in the paper. It contains:
  1. (1)
    A pattern-based analysis of the 2024 workshops, using the patterns from PoPII, and exploring the extent to which these patterns were used (implicitly or explicitly). A summary appears in Table LABEL:workshop-summary in Section 3.2.
  2. (2)
    Potential new PLACARD patterns, mentioned in the paper.
  3. (3)
    Candidate Patterns for Pattern-competent agents.
  4. (4)
    Candidate Patterns for Pattern-competent Agent Environments.
Pattern analysis of the 2024 workshops
One immediate result was the identification of of a proto-pattern, “Have Multiple Leaders Onsite,” addressing challenges in hybrid-facilitated events:
Context: Interactive conference workshops with both in-person and remote participants.
Problem: A single onsite facilitator struggles to manage breakout groups, technology, and overall flow, leading to gaps in support and reduced engagement (especially for remote attendees).
Solution: Deploy multiple co-facilitators physically onsite to coordinate activities, handle logistics, and engage with participants. Onsite leaders can divide roles (e.g., one per breakout table or task), ensuring better guidance and responsiveness, while remote facilitators focus on online participants.
A. Identifying Themes
DÉRIVE COMIX
Application: Participants used cards (object, value, body/territory, situation) as interpretive prompts to spark imagination—mirroring urban dérive in a conceptual sense. Observation: The card game facilitated divergent thinking akin to exploring unfamiliar terrain.
CONTEXT SETTING
Application: Early slides and explanations introduced key concepts (e.g., design patterns, CLA). Opportunity: Participant feedback highlighted a need for clearer upfront orientation. Improved alignment with this pattern could increase accessibility.
DO YOUR RESEARCH
Application: Past workshop experiences were integrated into the design; earlier uses of CLA and PAR informed this session. Example: Reference to Anticipation 2019 and prior events illustrates “pre-research” informing present design.
B. Organizing Structure
GOING META
Application: The project itself exemplifies recursive methodology—using PLACARD to refine PLACARD. Highlight: Participants were prompted to reflect on and imagine both the methodology and the subject (AI futures).
PATTERN LANGUAGE COMPONENTS
Partial: The workshop design showcases layering (CLA + DPL + PAR) but participants struggled to parse the conceptual load quickly. Recommendation: Use lighter keywords (HOWEVER, BECAUSE, SPECIFICALLY) to scaffold understanding incrementally.
FACILITATOR ROLES / FUNCTIONAL ROLES
Application: Hybrid facilitation with distinct roles (onsite leader, remote supporters) models distributed facilitation. Opportunity: Could be made more explicit—e.g., naming “Linker,” “Reflector,” or “Wrinkler” roles to participants.
C. Making it Actionable
THE FUTURE BEGINS NOW
Application: The card game generated imagined futures that participants shared and developed. Gap: While creative ideas emerged, follow-through actions or small-scale pilots were not described. An actionable step could have enhanced continuity.
PROJECT ACTION REVIEW (PAR)
Application: Strong use of the PAR pattern at the end of the session to reflect, share learnings, and gather insights. Insight: Participants gained confidence and intended to reuse techniques—an indicator of success.
REINFUSE EXPERTISE
Application: Participants brought diverse professional and cultural perspectives. Expertise was implicitly reintegrated via the shared reflection on AI futures and governance. Suggestion: A more explicit arc—temporarily setting aside expertise to foster common ground—could enrich this pattern’s application.
Potential new PLACARD patterns
@@ A. IDENTIFYING THEMES
Context: You are facilitating a workshop intended to generate new ideas or shared understanding. If: You want to activate creativity and presence, But: participants arrive cold or disconnected from each other and the topic, Then: Begin with a sensory or embodied activity (e.g., rhythm, image-based storytelling, gesture) that creates shared energy and primes pattern recognition. Evidence base from related literature: Techniques such as ”story circles” (participatory arts), embodied methods in design ethnography, and rhythm-based facilitation in music education all affirm the value of grounding group attention through shared sensory engagement. These strategies mirror the function of PAR alternatives that aim to surface and align affective states as a precondition for inquiry.
@@
Context: You are introducing participants to a method that is unfamiliar or complex (like PLACARD). If: You want them to participate actively, But: the framing is opaque or abstract, Then: Explicitly describe the method, its purpose, and how today’s session fits into the broader picture. Evidence base from related literature: This mirrors practices from facilitation toolkits like Liberating Structures (e.g., ”Purpose-to-Practice”) and the role of explicit framing in adult learning theory, where naming the structure of engagement is key to building learner autonomy and motivation.
@@
Context: You are using prompts (cards, challenges, scenarios) to guide group inquiry. If: You offer a situation that feels flat or generic, But: participants want something meaningful to anchor their play, Then: Invite participants to select, remix, or define their own situation to explore. Evidence base from related literature: User-centered design emphasizes co-defining the design challenge. Futures methods such as the ”Futures Triangle” and participatory foresight often encourage the co-creation of situational prompts to enhance contextual fit and participant investment.
@@ B. Organizing Structure
Context: You are working with participants who may be new to design patterns. If: You want people to identify and name patterns, But: the idea of a ”pattern” is still abstract, Then: Seed early examples or light prompts (”What did you do just now that worked?”) to build fluency. Evidence base from related literature: This aligns with double-loop learning (Argyris and Schön), where patterns of thought and action are surfaced through inquiry. It’s also reinforced in cognitive apprenticeship, which models strategies before expecting learners to apply them.

### v2

@@ 7.1. Limitations and Lessons Learned
Although the outcomes of our workshops — both in-person and simulated — were promising, several limitations emerged that highlight areas for refinement, both in the method itself and in our implementation of it. Generally, these limitations taught us lessons about blocks that could be overcome by applying appropriate precautions in future work.
…in in-person workshops
@@
Cognitive demands of CLA+PAR+DPL+other activities: The abstraction and depth of Causal Layered Analysis (CLA) alone might have been enough to overwhelm participants, and often the topics we were presenting were considerably more complex. Provide lightweight onboarding tools, and simplify the workshop, to ease entry into layered/multiplex analysis.
Logistical overhead of documentation: Without dedicated roles or tooling, PLACARD’s original emphasis on generating shared text in various formats was burdensome. Support tool development and role formalization (e.g., for pattern capture). Design light-touch alternatives for low-resource contexts.
Short workshop windows (60–90 mins): Limited time forced trade-offs, with later phases (e.g., PAR) often rushed or skipped. Recognize Timebox the Core as a necessary pattern—but pair it with mechanisms for asynchronous reflection or follow-up synthesis (with or without the workshop participants)
@@
Structured methods sometimes crowded out emotional or disruptive forms of insight. Integrate methods or prompts that scaffold expressive, intuitive, and affective contributions—e.g., through metaphor, narrative, or art-based reflection.
Participants come with prior expectations around the topic and format (e.g., AI, card games) We need to do some bridge work to negotiate between participants’ expectations/imagination and the workshop activities; ideally this happens with together, not after-the-fact.
…in the virtual workshop
Limitation Lessons Learned
Simulated agents were archetypal, not adaptive: The multi-agent models lacked real-time learning or memory; they played out scripted reasoning paths. Treat the current set of simulations as design probes rather than predictive models. Use them to explore a design space and pattern viability, not for behavioral realism.
@@
No persistent identity or memory across sessions: Agents did not evolve across workshops, limiting insight into long-term learning. ChatGPT does not run a loggable agent simulation framework, and it would need to be coaxed to produce plausible transcripts from even one virtual workshop. Incorporate persistent memory structures and episodic traces to study cumulative reflection, pattern evolution, and institutionalization.
Language model constraints shaped outcomes: Agent behaviors emerged from a coherence-biased generative system, not open-ended cognition. The simulations must be interpreted with epistemic humility. Their role is speculative, not definitive—ideal for surfacing design tensions, not modeling reality.
The experiments were somewhat hard to control (e.g., proposing to evolve patterns that were not part of the input set); while the methods are repeatable, the exact outputs are not replicable. The next round of experiments along these lines would more robust if we used a more structured purpose-built MAS.
@@ 7.2. Related Work
Pitt et al. (2012) explore how institutions can self-organize by formalizing socio-economic principles that guide agent behavior within Multi-Agent Systems (MAS). Drawing inspiration from Elinor Ostrom’s work, they propose an axiomatic design approach that captures principles such as proportionality, accountability, and collective choice in a machine-interpretable form. Where Pitt et al. emphasize axioms as foundational building blocks, our approach leverages design patterns as intermediate representations that bridge human practice and machine reasoning.  In a more recent paper, Mertzani and Pitt (2024) present a methodological framework for integrating power sensitivity into the design of cyber-physical systems by drawing on second- and third-order cybernetics. They propose pattern-based design methods that foreground the observer’s position, across different layers of embeddedness—with beneficial applications of technology ranging from “Saves Physical and Cognitive Resources” to “Solves Complex Problems” to “Deals with Social Constructions” to “Mitigates Existential Crises”.
7.3. Summary
We introduced the pattern Patterns Embody Intelligence before embarking on the more contextual (environmental, developmental) thinking decribed above. That thinking can now be summarised as an extension and reframing of the pattern. This adaptation makes the edges between the sensory/cognitive/motor aspects less sharp, and the pattern more explicitly engages the context.
The most subtantive change, however, is that we are now in a position to cross-references evidence (and potential implementation ideas) from our pattern-capable agent patterns.
Patterns Embody Intelligence
x
Context: Intelligence does not arise in isolation. (Shared Memory Architecture: “a persistent shared memory, enabling learning and reuse”)
—Perception, understanding, and action interweave (Reflection Scaffold Upgrade: “developmental learning depends on scaffolding”).
If: You employ a sensory method to foreground a sub-context, a cognitive method to explore its tensions and meanings, and a motor method to act within it (Candidate Move Generation: “Generate a set of plausible candidate actions”; Pattern Activation: “latent affordances become real decisions”).
However: Even so, your efforts may falter, and your methods may prove ineffective. (Tension Detection: “Flag the tension explicitly”).
—Insights may remain abstract, brittle, or out of sync with lived conditions. (Missing Architecture Agent: “not yet participating in the ecology of practice”).
Then: Hone your approach by identifying new forms of intelligence—e.g., collective, embodied, epistemic, imaginative, even ‘artificial’ (Role Reveal: “successful coordination depends on transparency of roles”)
—and engage them for purposes of methodological improvement (Timebox the Core: “prioritize core pattern loops”).
Because: Patterns are more than just methods: they are, themselves, scaffolds of intelligence. Evolving them reshapes the systems they act within, even if they are, for now, mediated primarily by words (Feedback Rhythms: “tune system-wide epistemic strategies”).
@@ 8. Conclusion
Looking back over the limitations and lessons learned, it’s worth drawing out one top-level remark. ChatGPT is adept at text processing, and we have put it to use significantly in our work on this paper. It was able to generate and discuss design patterns that describe pattern-capable agents, and to mock up virtual multi-agent workshops at different levels of detail. In particular, our second virtual workshop uses a dialogue-based protocol to simulate a discussion between pattern-capable agents. We hope this paper might inspire readers to explore potential combinations of LLMs and design patterns—but we see that as a landmark rather than a destination.
We have previously explored design fiction as a way of working (by hand (Danoff et al., 2019)), and we are aware that fictional works can have unexpected real-world ramifications, e.g., (Wikipedia contributors, 2025). In this case, we’ve demonstrated that a popular contemporary AI system, is, itself, “pattern-capable’.
Including a computational use case within the paper suggested a key amendment to PLACARD: adding a fourth environment dimension to the pattern catalogue. To close this paper, we put forward the following CLA, which brings together the “human” and “AI” use cases and outlines a trajectory of future work. In the context of this paper, the CLA layers can be aligned with the expanded view of Patterns Embody Intelligence, in which patterns encompass sensory (litany), cognitive (system), motor (worldview), and environment (myth) dimensions.
@@
In a well-know address an audience of computer scientists, Alexander (1999) said: “I don’t really see discussion about what, collectively, computer scientists are supposed to be doing with all these programs. How are they supposed to help the Earth?” That discussion has of course been going on in the pages of journals like AI & Society. However, the links to more technical discussion of patterns (for example, in works like (Elish and Hwang, 2016; Heiland et al., 2023)) are less obvious. We suggest that both the terms and stakes of the discussion have, indeed, changed. Connections need to be drawn between visionary thinking, tried-and-tested design approaches, and the practical, visible, and widely discussed elements of current AI discourse, including the promise and perils of large language models (LLMs) and the early success of LLM-associated code-writing agents.
@@
As a next step, we propose pattern-based epistemic scaffolding for multi-agent collaboration. (In the current paper, this has been prototyped at the design fiction level.)
@@
Agents trained to recognize and apply design patterns in their interactions. They exhibit situated competence and can engage in structured reasoning, scenario exploration, and inquiry.
@@
These agents also look out for opportunities to revise existing patterns and propose new ones, participating in iterative design and reflective cycles. They enable epistemic learning at the system level, coordinating multi-agent behavior across contexts.
@@
Inspired by Elinor Ostrom’s principles for governing commons, we treat the MAS as a social-technological institution, or federation of institutions, that evolves through interaction. (Again, we have presented a very limited early text-based prototype.) Bridges between more technical and more policy-oriented pattern languages (such as the pattern language for developing regions of Mehaffy et al. (2020)) should be pursued.
@@
This layer operationalizes developmental patterns that support reflexivity, feedback, nested structures, and shared norm evolution across MAS environments. These agents co-design systems of engagement, monitor norm congruence, and resolve governance disputes. They instantiate a worldview in which intelligence is systemic, collective, and emergent.
@@
Beyond technical systems, our proposal invites reflection on how other systems—academic, policy, scientific—can themselves become adaptive, transparent, and pattern-literate.
@@
Humans can use software tools to model, simulate and improve human institutions—e.g., Farmer (2024) focuses on markets and touches more broadly on how “agent-based models could be used to guide us through some of our biggest problems.” By working with agents that are both pattern- and institution-capable, we would gain assurances not only about avoiding “run-away AI” scenarios, but also about avoiding other tragedies of the commons. Alexander asked, “How are programs supposed to help the Earth?” Our answer is: not merely by embedding patterns in code, or even by building pattern-based programs, but by using patterns to shape and debug the programming of institutions, infrastructures, and our shared planetary futures.
@@ 9. Acknowledgements
We thank our PLoP paper shepherd Valentio Vranic for contructive comments and feedback on our submitted paper and its revisions. We also wish to acknowledge discussions with Cris Viseu around our 2025 EuroPLoP workshop. As mentioned, this paper also benefited from extensive interaction with OpenAI’s ChatGPT, which we used to support the analysis in Section 4, to run the virtual workshop in Section 6, to author the patterns in the appendix. We used both ChatGPT and a range of similar tools for assistance with early-stage ideation, outlining, and drafting individual sections. [^5] https://groups.google.com/g/peeragogy/c/XM7RKMDdcT4 As indicated in the introduction, the contents of the paper have been thoroughly revised since then.
@@
This appendix supports the analysis in the paper. All patterns were developed by ChatGPT in response to our prompting, [^6] E.g., “Are there new patterns that we can add that would extrapolate what we’ve learned from the workshops?” […] “For each of the patterns that you proposed, could you add a new subsection on ‘evidence base from related literature’, referencing the analogues to PAR, CLA, and DPL that you found before?” sometimes with minor changes. The appendix contains: (1) potential new patterns for structuring in-person workshops; (2) potential new patterns for pattern-competent agents; (3) potential new patterns for pattern-competent agent environments; (4) a small number of new or revised patterns from the simulated workshop in Section 6.2; (5) further supplementary pattern(s) that inform our the discussion but which were not used in the experiments.
(1) Potential new patterns for structuring workshops
These were generated following the analysis of in-person workshops in terms of their use of PAR, CLA, and DPL and the existing PLACARD patterns (Section 4). The patterns were subsequently made available to virtual workshop participants, along with the PLACARD pattern language, in the experiment discussed in Section 6.1.
@@ A. IDENTIFYING THEMES
Context: You are facilitating a workshop intended to generate new ideas or shared understanding. If: You want to activate creativity and presence, But: participants arrive cold or disconnected from each other and the topic, Then: Begin with a sensory or embodied activity (e.g., rhythm, image-based storytelling, gesture) that creates shared energy and primes pattern recognition. Evidence base from related literature: Techniques such as “story circles” (participatory arts), embodied methods in design ethnography, and rhythm-based facilitation in music education all affirm the value of grounding group attention through shared sensory engagement. These strategies mirror the function of PAR alternatives that aim to surface and align affective states as a precondition for inquiry.
@@
Context: You are introducing participants to a method that is unfamiliar or complex (like PLACARD). If: You want them to participate actively, But: the framing is opaque or abstract, Then: Explicitly describe the method, its purpose, and how today’s session fits into the broader picture. Evidence base from related literature: This mirrors practices from facilitation toolkits like Liberating Structures (e.g., “Purpose-to-Practice”) and the role of explicit framing in adult learning theory, where naming the structure of engagement is key to building learner autonomy and motivation.
@@
Context: You are using prompts (cards, challenges, scenarios) to guide group inquiry. If: You offer a situation that feels flat or generic, But: participants want something meaningful to anchor their play, Then: Invite participants to select, remix, or define their own situation to explore. Evidence base from related literature: User-centered design emphasizes co-defining the design challenge. Futures methods such as the “Futures Triangle” and participatory foresight often encourage the co-creation of situational prompts to enhance contextual fit and participant investment.
@@ B. Organizing Structure
Context: You are working with participants who may be new to design patterns. If: You want people to identify and name patterns, But: the idea of a “pattern” is still abstract, Then: Seed early examples or light prompts (“What did you do just now that worked?”) to build fluency. Evidence base from related literature: This aligns with double-loop learning (Argyris and Schön), where patterns of thought and action are surfaced through inquiry. It’s also reinforced in cognitive apprenticeship, which models strategies before expecting learners to apply them.